Circular No. 7368

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1999 WJ_{7} (KORLEVIĆ) An apparently asteroidal object discovered by Korado Korlević at

Višnjan with a 0.41-m f/4.3 reflector + CCD (the discovery observation below was first published on MPS 7912) was indicated on some of his December images to be possibly “fuzzy”, and cometary nature has been confirmed by C. Hergenrother and S. Larson, Lunar and Planetary Laboratory, who found a 13′′ coma elongated in p.a. 80^{o} on a 600-s co-added R-band exposure taken on 2000 Feb. 7.25 UT with the 1.54-m Catalina reflector.

1999 UT  α_{2000}  δ_{2000}  m_{2}
Nov. 28.94427  4^{h}34^{m}56.^{s}58  +24^{o}51′51′′.8  17.9

Additional observations and orbital elements (T = 2000 Feb. 15.2 TT, q = 3.17 AU, i = 2^{o}.98, P = 9.97 yr) appear on MPEC 2000-D08.

COMET C/2000 B4 (LINEAR) Another apparently asteroidal object was reported by LINEAR and

posted on the NEO Confirmation Page. This object has the orbit of a centaur and was noted as appearing perhaps slightly diffuse (P. Kušnirak, Ondřejov, 0.65-m reflector, Feb. 10) and ‘soft’ and slightly larger than star images (D. Balam, Victoria, 1.82-m reflector, Feb. 11). Astrometry and orbital elements (T = 2000 June 24.6 TT, q = 6.82 AU, i = 15^{o}.9, P = 77 yr) appear on MPEC 2000-D09.

2000 UT  α_{2000}  δ_{2000}  m_{2}
Jan. 29.24832  8^{h}08^{m}15.^{s}43  +37^{o}03′04′′.0  19.4

COMET C/2000 CT_{54} (LINEAR) Yet another apparently asteroidal LINEAR object that was posted on

the NEO Confirmation Page was noted to have a 15′′–16′′ tail toward the north-northwest on Feb. 12 by J. G. Ries, McDonald Observatory (0.76-m reflector). Astrometry and parabolic orbital elements (T = 2001 June 18.6 TT, q = 3.13 AU, i = 49^{o}.2) appear on MPEC 2000-D10.

2000 UT  α_{2000}  δ_{2000}  m_{2}
Feb. 2.44006  14^{h}03^{m}40.^{s}19  −21^{o}20′28′′.4  18.9
2000 February 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green